OpenStudy (jetman24):

Which equation represents exponential growth and has a vertical stretch of 2? A. y = 2(3x) B. y = 3(.11)x C. y = 2(1/2)x D. y = x(1/6)2

OpenStudy (giantrobot):

Can you send a screenshot? I think you misquoted it, because right now the equations look suspicious for the kind of question you're asking.

OpenStudy (jetman24):

how do I send a screenshot? I cant paste it into open study.

OpenStudy (giantrobot):

.....I would suggest Gyazo, but how about using the equation function to copy the qeuestion exactly?

OpenStudy (jetman24):

I apologize that I am so stupid right now but how do I do that? Because what I did was just copy and paste the question from my school website (I do online school).

OpenStudy (giantrobot):

Take a screenshot, save in paint and attach it to the open study.

OpenStudy (jetman24):

there I attached it

OpenStudy (giantrobot):

Oh yesssss. Lets eliminate the answers one at a time. First we start with a definition of exponential growth. Accelerating growth. Now we look at the answers. We can easily remove D because it can be written down as y=kx, nothing exponential here, it's linear with a high start. So for A, B and C, B and C are eliminated because the number attached to the x is less than one, it may seem counter-intuitive, but that doesn't cause exponential growth leaving A. Now why don't B and C have exponential growth?\[\frac{ 1 }{ 2 }\times \frac{ 1 }{ 2 } \] can be written as \[\frac{ 1 }{ 2 }^{2}\] Which gets \[\frac{ 1 }{ 4}\] A smaller number, so yeah. A is answer.

OpenStudy (jetman24):

Thank you for the help I think I gave you a medal by clicking best answer.